• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nilanga

    Log entry time 08:00:52 on March 9,2005

    Entry number 140942

    keyword=Shift Summary

    3.15 Beam in hall, start harp scans, beam center etc
    beam size 87 um x 112 um

    4.00 Ramped up the Bigbite current to 518 A, while MCC monitored the beam on dump. No change in beam was observed.

    4.10 Moved to no target; run with 5 uA

    4.20 Moved to slanted 12C target

    5.30 While data being analyzed for the gain matching, a test run on LH2

    6.30 Moving to momentum scan while waiting for analysis: 5 uA gives 250kHz on left arm; take only 3 uA, or 2 uA
    6.35 Start momentum scan


    Run # target PS3 PS6
    2470 empty 1 2
    2471 sla. C 120 500
    2472 junk
    2473 4cm LH2 1000 20

    Momentum scan:

    Run # angle P0 target current
    2474 15 1.137 LH2 3 uA
    2475 15 LH2 2 uA
    2476 15 sla.12C 2 uA
    2477 15 sla.12C 2 uA